OpenStudy (anonymous):

I need help. An open tank contains a layer of oil floating on top of a layer of water (of density 1000 kg/m^3) that is 3.0 m thick, as shown. What must be the thickness of the oil layer if the gauge pressure at the bottom of the tank is to be 6.4×10^4 Pa? The density of the oil is 820 kg/m^3.

OpenStudy (anonymous):

The gauge pressure at the bottom of the tank is simply equal to the sum of the pressure exerted by the oil and water so; \[P =(\rho _{water}*g*h _{_{water}}) + (\rho _{oil}*g*h _{oil})\] where g =9.81 m/s^2 then you can substitute and solve it easily

OpenStudy (anonymous):

You have also to add the pressure exerted by the air because its an open tank. Sorry i forgot it

OpenStudy (anonymous):

This is the information I typed up 6.4x10^4 = (1000 x 9.8 x 3) + (820 x9.8 x h) H= 63996 Which did not make sense why it's so thick How do I add the pressure exerted by air Thank you

OpenStudy (anonymous):

Ops! I think adding the air pressure is wrong because they gave you already the gauge pressure not the absolute. If you substitute you must get h equal to 4.3m \[h= [(6.4*10^4)-(1000 * 9.8 * 3)]/(820 * 9.8) = 4.3 m\]

OpenStudy (anonymous):

Great, thank you